# Confidence-driven Sampling for Backdoor Attacks

### v1

@@ Abstract
Backdoor attacks aim to surreptitiously insert malicious triggers into DNN models, granting unauthorized control during testing scenarios. Existing methods lack robustness against defense strategies and predominantly focus on enhancing trigger stealthiness while randomly selecting poisoned samples. Our research highlights the overlooked drawbacks of random sampling, which make that attack detectable and defensible. The core idea of this paper is to strategically poison samples near the model’s decision boundary and increase defense difficulty. We introduce a straightforward yet highly effective sampling methodology that leverages confidence scores. Specifically, it selects samples with lower confidence scores, significantly increasing the challenge for defenders in identifying and countering these attacks. Importantly, our method operates independently of existing trigger designs, providing versatility and compatibility with various backdoor attack techniques. We substantiate the effectiveness of our approach through a comprehensive set of empirical experiments, demonstrating its potential to significantly enhance resilience against backdoor attacks in DNNs.
@@ 1 Introduction
During DNN training on large datasets or third-party collaborations, there exist concerns about potential malicious triggers injected into the model. These intrusions can lead to unauthorized manipulation of the model’s outputs during testing, causing what is commonly referred to as a “backdoor” attack (Li et al., 2022; Doan et al., 2021a). To elaborate, attackers can inject triggers into a small portion of training data in a specific manner. Attackers may then provide either the poisoned training data or backdoored models trained on it to third-party users, depending on their capabilities (Li et al., 2022). In the inference stage, the injected backdoors are activated via triggers, causing triggered inputs to be misclassified as a target label. To date, numerous backdoor attack methods, such as BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), label-consistent (Turner et al., 2019), have demonstrated strong attack performance. These methods consistently achieve high attack success rates while maintaining a high accuracy on clean data within mainstream DNNs.
An important research direction in backdoor attacks is to enhance of the stealthiness of poisoned samples while concurrently ensuring their effectiveness. Most efforts in this research line have been made to trigger design (e.g., hidden triggers Saha et al., 2020, clean-label (Turner et al., 2019)). However, in the vast majority of existing attack methods, samples are randomly chosen from the clean training set for poisoning. However, our preliminary study (in Section 4.1) observes that the “random sampling” strategy is highly related to the possibility of the poisoning samples to be detected by the defenders. Moreover, it is totally practical and feasible for the attackers to choose the poisoning samples from the training set with preference. In fact, it is a common setting to assume that the attacker has the knowledge of the victim model’s training dataset for sample selection. For example, the victim models can be trained on downloaded online datasets, which are provided by the attacker (Li et al., 2022). Similarly, the attacker can also act as model providers to directly provide the backdoored models (Nguyen & Tran, 2021; Doan et al., 2021b). Therefore, there is plenty of room for discussing the sampling strategies in this scenario, which raises the question: Is there a better sampling strategy to improve the stealthiness of backdoors?
@@
To answer this question, in Section 4.1, we first take a closer look at the random sampling strategy, by investigating the latent space of the backdoored model. From the visualizations in Figure 1, we draw two interesting findings: First, most of the randomly chosen samples are close to the center of their true classes in the latent space; second, the closer a sample is from its true class on the clean model, the further it gets from the target class on the backdoored model. These two observations reveal important clues about the “stealthness” of the random sampling strategy, which suggest the randomly sampled data points may cause them to be easier to be detected as outliers. To have a deeper understanding, we further build a theoretical analysis of SVM in the latent space (Section 4.3) to demonstrate the relation between the random sampling strategy and attack stealthiness. Moreover, our observations suggest an alternative to the random sampling—it is better to select samples that are closer to the decision boundary. Our preliminary studies show that these boundary samples can be manipulated to be closer to the clean samples from the target class, and can greatly enhance their stealthness under potenital outlier detections (see Figure 0(c) and 0(d)).
@@
Inspired by these discoveries, we propose a novel method called confidence-driven boundary sampling (CBS). To be more specific, we identify boundary samples with low confidence scores based on a surrogate model trained on the clean training set. Intuitively, samples with lower confidence scores are closer to the boundary between their own class and the target class in the latent space (Karimi et al., 2019) and can avoid vulnerabilities brought by random samplings. Therefore, this strategy makes it more challenging to detect attacks. Moreover, our sampling strategy is independent from existing attacking approaches which makes it exceptionally versatile. It easily integrates with a variety of backdoor attacks, offering researchers and practitioners a powerful tool to enhance the stealthiness of backdoor attacks without requiring extensive modifications to their existing methods or frameworks. Extensive experiments combining proposed confidence-based boundary sampling with various backdoor attacks illustrate the advantage of the proposed method compared with random sampling.
@@ 2.1 Backdoor attacks and defenses
As mentioned in the introduction, backdoor attacks are shown to be a serious threat to deep neural networks. BadNet (Gu et al., 2017) is the first exploration that attaches a small patch to samples and therefore introduces backdoors into a DNN model. After that many efforts are put into developing advanced attacks to either boost the performance or improve the resistance against potential defenses. Various trigger designs are proposed, including image blending (Chen et al., 2017), image warpping (Nguyen & Tran, 2021), invisible triggers (Li et al., 2020; Saha et al., 2020; Doan et al., 2021b), clean-label attacks (Turner et al., 2019; Saha et al., 2020), sample-specific triggers (Li et al., 2021b; Souri et al., 2022), etc. These attacking methods have demonstrated strong attack performance (Wu et al., 2022). In the meanwhile, the study of effective defenses against these attacks also remains active. One popular type of defense depends on detecting outliers in the latent space (Tran et al., 2018; Chen et al., 2018; Hayase et al., 2021; Gao et al., 2019; Chen et al., 2018). Other defenses incorporate neuron pruning (Wang et al., 2019), detecting abnormal labels (Li et al., 2021a), model pruning (Liu et al., 2018), fine-tuing (Sha et al., 2022), etc.
@@ 2.2 Samplings in backdoor attacks
While the development of triggers in backdoor attacks attracts much attention, the impact of poisoned sample selections is rarely explored. As far as we know, Xia et al. (2022) is the only work focusing on the sampling method in backdoor attacks. They proposed a filtering-and-updating strategy (FUS) to select samples with higher contributions to the injection of backdoors by computing the forgetting event (Toneva et al., 2018) of each sample. For each iteration, poison samples with low forgetting events will be removed and new samples will be randomly sampled to fill up the poisoned training set. Though this method shows improvement in performance, it ignores the backdoor’s ability to resist defenses, known as the ’stealthiness’ of backdoors. To the best of our knowledge, we are the first work study the stealthiness problem from the perspective of sampling.
@@ 3 Definition and Notation
In this section, we will introduce preliminaries about backdoor attacks, including the threat model discussed in this paper and a general pipeline that is applicable to many backdoor attacks.
@@ 3.1 Threat model
We follow the commonly used threat model for the backdoor attacks (Gu et al., 2017; Doan et al., 2021b). We assume that the attacker has access to the clean training set and can modify any subset from the training data. Then the victim trains his own models on this data and the attacker has no knowledge of this training procedure. In a real-world situation, attackers can upload their datasets to the internet. They can sneakily insert backdoors into their data and then share it with victims, who unknowingly use it to train their own models (Gu et al., 2017; Chen et al., 2017). Note that many existing backdoor attacks (Nguyen & Tran, 2021; Turner et al., 2019; Saha et al., 2020) already adopt this assumption and our proposed method does not demand additional capabilities from attackers beyond what is already assumed in the context of existing attack scenarios.
@@ 3.2 A general pipeline for backdoor attacks
(2) Trigger injection. Attackers design some strategies $T$ to inject the trigger $t$ into samples selected in the first step. In specific, given a subset of data $U$, attackers generate a poisoned set $T(U)$ as:
@@
where $G_{t}(x)$ is the attacker-specified poisoned image generator with trigger pattern $t$ and $S$ indicates the attacker-specified target label generator. After training the backdoored model $f(\cdot;\theta^{b})$ on the poisoned set, the injected backdoor will be activated by trigger $t$. For any given clean test set $D_{te}$ , the accuracy of $f(\cdot;\theta^{b})$ evaluated on trigger-embedded dataset $T(D_{te})$ is referred to as success rate, and attackers expect to see high success rate on any clean samples with triggers embedded.
@@ 4 Method
In this section, we will first analyze the commonly used random samplings, and then introduce our propose method as well as some theoretical understandings.
@@ 4.1 Revisit random sampling
Visualization of Stealthiness. Random sampling selects samples to be poisoned from the clean training set with the same probability and is commonly used in existing attacking methods. However, we suspect that such unconstrained random sampling is easy to be detected as outliers of the target class in the latent space. To take a look at the sample distribution in latent space, we first conduct TSNE (Van der Maaten & Hinton, 2008) visualizations of clean samples from the target class, and the poisoned samples which could be originally from other class but labeled as the target class. We consider these poisoned samples are obtained by two representative attack algorithms, BadNet (Gu et al., 2017) and Blend (Chen et al., 2017) both of which apply random sampling, on CIFAR10 (Krizhevsky et al., 2009), in Figure 0(a) and 0(b). In detail, the visualizations show the latent representations of samples from the target class, and the colors red and blue indicate poisoned and clean samples respectively. It is obvious that there exists a clear gap between poisoned and clean samples. For both attacks, most of the poisoned samples form a distinct cluster outside the clean samples. This will result in separations in latent space which can be easily detected by possible defenses. For example, Spectral Siginiture (Tran et al., 2018), SPECTRE (Hayase et al., 2021), SCAn (Tang et al., 2021) are representative defenses relying on detecting outliers in the latent space and show great power defending various backdoor attcks (Wu et al., 2022).
@@
Relation between Stealthiness & Random Sampling. In our study, we also observe the potential relation between random sampling and the stealthiness of backdoors. To elaborate, we further calculate the distance from each selected sample (without trigger) to the center [^1] The average of sample representations within this class. of their true classes computed on the clean model, which is denoted as $d_{o}$. As seen in Figure 1(a) and 1(b), random sampling tends to favor samples that are close to their true classes. However, we find $d_{o}$ may have an obvious correlation with the distance between the sample and the target class which we visualize in the previous Figure 1. Formally, we define the distance between each selected sample (with trigger) and the center of the target class computed on the backdoored model, as $d_{t}$. From Figure 1(c) and 1(d), we observe a negative correlation between $d_{t}$ and $d_{o}$, indicating that samples closer to the center of their true classes in the clean model tend to be farther from the target class after poisoning and thus easier to detect. These findings imply that random sampling often results in the selection of samples with weaker stealthiness. Our observations also suggest that samples closer to the boundary may lead to better stealthiness, and motivate our proposed method.
@@ 4.2 Confidence-driven boundary sampling (CBS)
To solve this issue, we consider the confidence score. To be more specific, we inherit the notations from Section 3.2 and further assume there exist $K$ classes, i.e. $\mathcal{Y}=\{1,...,K\}$, for simplicity. Let $f(\cdot;\theta)$ denote a classifier with model parameter $\theta$, and the output of its last layer is a vector $z\in\mathbb{R}^{K}$. Confidence score is calculated by applying the softmax function on the vector $z$, i.e. $\bm{s_{c}(f(x;\theta))=\sigma(z)}\in[0,1]^{K}$, where $\sigma(\cdot)$ is the softmax function. This confidence score is considered the most accessible uncertainty estimate for deep neural network (Pearce et al., 2021), and is shown to be closely related to the decision boundary (Li et al., 2018; Fawzi et al., 2018). Since our primary goal is to identify samples that are closer to the decision boundary, we anticipate finding samples with similar confidence for both the true class [^2] For a correctly classified sample, the true class possesses the largest score. and the target class. Thus, we can define boundary samples as:
@@
$|s_{c}(f(x;\theta))_{y}-s_{c}(f(x;\theta))_{y^{\prime}}|\leq\epsilon$ (2)
@@
The proposed Confidence-driven boundary sampling (CBS) method is based on Definition4.1. In general, CBS selects boundary samples in Definition 4.1 for a given threshold $\epsilon$. Since we assume the attacker has no knowledge of the victim’s model, we apply a surrogate model like what black-box adversarial attacks often do (Chakraborty et al., 2018). In detail, a pre-trained surrogate model $f(\cdot;\theta)$ is leveraged to estimate confidence scores for each sample, and $\epsilon$-boundary samples with pre-specified target $y^{t}$ are selected for poisoning. The detailed algorithm is shown in Algorithm 1 in Appendix A.3. It is worth noting that the threshold $\epsilon$ is closely related to poison rate $p$ in Section 3.2, and we can determine $\epsilon$ based on $|U(y^{t},\epsilon)|=p\times|\mathcal{D}_{tr}|$. Since we claim that our sampling method can be easily adapted to various backdoor attacks, we provide an example that adapts our sampling methods to Blend (Chen et al., 2017), where we first select samples to be poisoned via Algorithm 1 and then blend these samples with the trigger pattern $t$ to generate the poisoned training set. Detailed algorithms can be found in Algorithm 2 in Appendix A.3.

### v2

# Stealthy Backdoor Attack via Confidence-driven Sampling

@@ Abstract
Backdoor attacks facilitate unauthorized control in the testing stage by carefully injecting harmful triggers during the training phase of deep neural networks. Previous works have focused on improving the stealthiness of the trigger while randomly selecting samples to attack. However, we find that random selection harms the stealthiness of the model. In this paper, we identify significant pitfalls of random sampling, which make the attacks more detectable and easier to defend against. To improve the stealthiness of existing attacks, we introduce a method of strategically poisoning samples near the model’s decision boundary, aiming to minimally alter the model’s behavior (decision boundary) before and after backdooring. Our main insight for detecting boundary samples is exploiting the confidence scores as a metric for being near the decision boundary and selecting those to poison (inject) the attack. The proposed approach makes it significantly harder for defenders to identify the attacks. Our method is versatile and independent of any specific trigger design. We provide theoretical insights and conduct extensive experiments to demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
While deep neural networks (DNNs) on large datasets and third-party collaborations demonstrate promising performance in various applications, concerns have been raised about potential malicious triggers injected into the models. These triggers lead to unauthorized manipulation of the model’s outputs during testing, causing a “backdoor” attack (Li et al., 2022; Doan et al., 2021a). In particular, attackers can inject triggers into a small portion of training data in a specific manner, then provide either the poisoned training data or backdoored models trained on it to third-party users (Li et al., 2022). In the inference stage, the injected backdoors are activated via triggers, causing triggered inputs to be misclassified as a target label. In the existing literature, many backdoor attack methods have been developed and demonstrate strong attack performance, e.g., BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), and label-consistent (Turner et al., 2019). These methods can achieve high attack success rates while maintaining a high accuracy on clean data within mainstream DNNs.
An important research direction in backdoor attacks is to enhance the stealthiness of poisoned samples while ensuring their effectiveness simultaneously. While trigger design (e.g., hidden triggers Saha et al., 2020, clean-label (Turner et al., 2019)) has been a primary focus in existing research, recent studies have increasingly explored sampling methods for selecting optimal data points for poisoning and trigger insertion. However, for sampling methods, most existing works (Toneva et al., 2018; Han et al., 2023; Li et al., 2023; Xia et al., 2023; Wu et al., 2023; Zhu et al., 2023) focus on the attacking effectiveness while ignoring the stealthiness of backdoors. Our preliminary study (in Section 4.1) observes that the randomly selected poisoned samples are highly likely to be detected by the defenders. Such a weakness raises a natural question:
Is there a better sampling strategy to enhance the stealthiness of backdoors?
To investigate this question, we follow the common understanding to assume that the attackers can access the training data while maybe only allowed to manipulate a part of the training data. For example, the attackers may contribute malicious data to publicly sourced datasets via uploading their own data online (Li et al., 2022). Besides improving the trigger pattern, they also need a sampling strategy to determine the data to update.
@@
To better understand the behavior of the backdoor attacks, in Section 4.1, we investigate the latent space of the backdoored model to take a closer look at the random sampling strategy. We draw two findings from the visualizations in Figure 1. First, most randomly chosen samples are close to the center of their true classes in the latent space. Second, the closer a sample is from its true class on the clean model, the farther it gets from the target class on the backdoored model (Section 4.1). These two observations reveal an important concern about the “stealthiness” of the random sampling strategy, where the randomly sampled data points may be easily detected as outliers. To gain a deeper understanding, we further build a theoretical analysis of SVM in the latent space (Section 4.3) to demonstrate the relation between the random sampling strategy and attack stealthiness. Moreover, our observations suggest an alternative to random sampling—it is better to select samples closer to the decision boundary. Our preliminary studies show that these boundary samples can be manipulated to be closer to the clean samples from the target class and can greatly enhance their stealthiness under potential outlier detection (see Figure 1(c) and 1(d)).
@@
Inspired by the above observations, we propose a novel method called confidence-driven boundary sampling (CBS). Specifically, we identify boundary samples with low confidence scores based on a surrogate model trained on the clean training set. Intuitively, samples with lower confidence scores are closer to the boundary between their own class and the target class in the latent space Karimi et al. (2019) compared to random samples. Therefore, this strategy makes it more challenging to detect attacks. Moreover, our sampling strategy is independent from existing attack approaches, making it exceptionally versatile. It can be easily integrated with various backdoor attacks, offering researchers and practitioners a powerful tool to enhance the stealthiness of backdoor attacks without requiring extensive modifications to their existing methods or frameworks. Extensive experiments combining proposed confidence-based boundary sampling with various backdoor attacks illustrate the advantage of the proposed method over random sampling.
@@ 2.1 Backdoor attacks and defenses
As mentioned in the introduction, backdoor attacks are shown to be a serious threat to DNN. BadNet (Gu et al., 2017) is the first exploration that attaches a small patch to samples to introduce backdoors into a DNN model. Later, many efforts are put into developing advanced attacks to boost the performance or improve the resistance against potential defenses. Various trigger designs are proposed, including image blending (Chen et al., 2017), image warpping (Nguyen & Tran, 2021), invisible triggers (Li et al., 2020; Saha et al., 2020; Doan et al., 2021b), clean-label attacks (Turner et al., 2019; Saha et al., 2020), sample-specific triggers (Li et al., 2021b; Souri et al., 2022), etc. These attacking methods have demonstrated strong attack performance (Wu et al., 2022). Meanwhile, the study of effective defenses against these attacks also remains active. One popular type of defense detects outliers in the latent space (Tran et al., 2018; Chen et al., 2018; Hayase et al., 2021; Gao et al., 2019; Chen et al., 2018). Other defenses incorporate neuron pruning (Wang et al., 2019), detecting abnormal labels (Li et al., 2021a), model pruning (Liu et al., 2018), fine-tuing (Sha et al., 2022), etc.
@@ 2.2 Samplings in backdoor attacks
Despite the development of triggers in backdoor attacks, the impact of poisoned sample selection is also attracting more and more attention. Xia et al. (2022) proposed a filtering-and-updating strategy (FUS) to select samples with higher contributions to the injection of backdoors by computing the forgetting event (Toneva et al., 2018) of each sample. For each iteration, poison samples with low forgetting events will be removed, and new samples will be randomly sampled to fill out the poisoned training set. Han et al. (2023); Li et al. (2023); Xia et al. (2023) followed this line and also adopted the forgetting score for sample selection. Wu et al. (2023); Zhu et al. (2023) leverages masks and $l_{2}$ distance in representation space respectively to improve the effectiveness of the backdoor. Though these works can improve the success rate of backdoor attacks via sample selection, they ignore the backdoor’s ability to resist defenses, known as the ’stealthiness’ of backdoors. To the best of our knowledge, we are the first to study the stealthiness problem from the sampling perspective.
@@ 3 Definition and Notation
This section introduces preliminaries about backdoor attacks, including the threat model considered in this paper and a general pipeline that is applicable to many attacks.
@@ 3.1 Threat model
We follow the commonly used threat model for the backdoor attacks (Gu et al., 2017; Doan et al., 2021b). We assume that the attacker can access the clean training set and modify a proportion of the training data. Then, the victim trains his own models on this data, and the attacker has no knowledge of this training procedure. In a real-world situation, the attacker can access some clean datasets, and modify a proportion of them by inserting triggers. Then they upload the poisoned data to the Internet and victims unknowingly download and use it for training (Gu et al., 2017; Chen et al., 2017). Note that many existing backdoor attacks (Nguyen & Tran, 2021; Turner et al., 2019; Saha et al., 2020) have already adopted this assumption, and our proposed method does not demand additional capabilities from attackers beyond what is already assumed in the context of existing attack scenarios. Furthermore, our method, detailed in Section 4, addresses practical scenarios where attackers are limited to poisoning samples from a specific subset, not the entire dataset, with empirical results in Section 5.5. For example, an attacker might only control their own data and not have access to alter public datasets.
@@ 3.2 A general pipeline for backdoor attacks
(2) Trigger injection. Attackers design a strategy $T$ to inject the trigger $t$ into samples selected in the first step. In specific, given a subset of data $U$, attackers generate a poisoned set $T(U)$ as:
@@
where $G_{t}:\mathcal{X}\rightarrow\mathcal{X}$ is the attacker-specified poisoned image generator with trigger pattern $t$, which satisfies the following constraints: $G_{t}(x)\neq x$ and $d(G_{t}(x),x)\leq\epsilon$ where $d$ is some distance function such as $l_{2}/l_{\infty}$ distance, and $S$ indicates the attacker-specified target label generator. After training the backdoored model $f(\cdot;\theta^{b})$, where $\theta^{b}$ denote parameters of the backdoored model, on the poisoned set, the injected backdoor will be activated by trigger $t$. For any given clean test set $D_{te}$, the accuracy of $f(\cdot;\theta^{b})$ evaluated on trigger-embedded dataset $T(D_{te})$ is referred to as success rate, and attackers also expect to see high accuracy on any clean samples without triggers.
@@ 4 Method
In this section, we will first analyze the commonly used random sampling method, and then introduce our proposed method as well as some theoretical understandings.
@@ 4.1 Revisit random sampling
Visualization of Stealthiness. Random sampling selects samples to be poisoned from the clean training set with the same probability and is commonly used in existing attacking methods. However, we suspect that such unconstrained random sampling is easy to detect as outliers of the target class in the latent space. To examine the sample distribution in the latent space, we first conduct TSNE (Van der Maaten & Hinton, 2008) visualizations for the backdoored model of (1) clean samples of the target class, and (2) the poisoned samples from other classes but labeled as the target class. We consider these poisoned samples are obtained by two representative attack algorithms, BadNet (Gu et al., 2017) and Blend (Chen et al., 2017) both of which apply random sampling, on CIFAR10 (Krizhevsky et al., 2009), in Figure 1(a) and 1(b). In detail, the visualizations show the latent representations of samples from the target class, and the colors red and blue indicate poisoned and clean samples respectively. One can see a clear gap between poisoned and clean samples. For both attacks, most of the poisoned samples form a distinct cluster outside the clean samples. This indicates a separation in latent space which can be easily detected by potential defenses. For example, Spectral Signature (Tran et al., 2018), SPECTRE (Hayase et al., 2021), SCAn (Tang et al., 2021) are representative defenses relying on detecting outliers in the latent space and show great power defending various backdoor attacks (Wu et al., 2022).
@@
Relation between Stealthiness & Random Sampling. In our study, we also observe the potential relation between random sampling and the stealthiness of backdoors 111 We provide a detailed discussion of “stealthiness” in Appendix 8.5.. To elaborate, we further calculate the distance from each selected sample (without trigger) to the center [^2] A formal definition of $d_{o}$ and $d_{t}$ is shown in Appendix 8.5 of their true classes computed on the clean model, which is denoted as $d_{o}$. As seen in Figure 2(a) and 2(b), random sampling is likely to select samples that are close to the center of their true classes. However, we find $d_{o}$ may have an obvious correlation with the distance between the sample and the target class which we visualize in the previous Figure 1. Formally, we define the distance between each selected sample (with trigger) and the center of the target class computed on the backdoored model as $d_{t}$.
From Figure 2(c) and 2(d), we observe a negative correlation between $d_{t}$ and $d_{o}$  [^3] We also include a discussion of the relationship between raw input space and latent space in Appendix 8.11., indicating that samples closer to the center of their true classes in the clean model tend to be farther from the target class after poisoning and thus easier to detect. These findings imply that random sampling often results in the selection of samples with weaker stealthiness. Our observations also suggest that samples closer to the boundary may lead to better stealthiness and motivate our proposed method.
@@ 4.2 Confidence-driven boundary sampling (CBS)
To solve this issue, we consider the confidence score. To be more specific, we follow the notations from Section 3.2 and further assume there exist $K$ classes, i.e., $\mathcal{Y}=\{1,...,K\}$, for simplicity. Let $f(\cdot;\theta)$ denote a classifier with model parameter $\theta$, and the output of its last layer is a vector $z\in\mathbb{R}^{K}$. Confidence score is calculated by applying the softmax function on the vector $z$, i.e. $\bm{s_{c}(f(x;\theta))=\sigma(z)}\in[0,1]^{K}$, where $\sigma(\cdot)$ is the softmax function.
This confidence score is considered the most accessible uncertainty estimate for deep neural network (Pearce et al., 2021) and is shown to be closely related to the decision boundary (Li et al., 2018; Fawzi et al., 2018). Since our primary goal is to identify samples that are closer to the decision boundary, we can find samples with similar confidence for both the true class [^4] For a correctly classified sample, the true class possesses the largest score. and the target class. Thus, we can define boundary samples:
@@
$|s_{c}(f(x;\theta))_{y}-s_{c}(f(x;\theta))_{y^{\prime}}|\leq\epsilon,$ (2)
@@
The proposed Confidence-driven boundary sampling (CBS) method is based on Definition 4.1. In general, CBS selects boundary samples in Definition 4.1 for a given threshold $\epsilon$. Since we assume the attacker has no knowledge of the victim’s model, we apply a surrogate model like what black-box adversarial attacks often do (Chakraborty et al., 2018). In detail, a pre-trained surrogate model $f(\cdot;\theta)$ is leveraged to estimate confidence scores for each sample, and $\epsilon$-boundary samples with pre-specified target $y^{t}$ are selected for poisoning. The detailed algorithm is shown in Algorithm 1 555 Discussion of computation overhead is included in Appendix 8.4. Note that the threshold $\epsilon$ is closely related to poison rate $p$ in Section 3.2, and we can determine $\epsilon$ so that $|U(y^{t},\epsilon)|=p\times|\mathcal{D}_{tr}|$. Since we claim that our sampling method can be easily adapted to various backdoor attacks, we provide an example that adapts our sampling methods to Blend (Chen et al., 2017), where we first select samples to be poisoned via Algorithm 1 and then blend these samples with the trigger pattern $t$ to generate the poisoned training set.
Algorithm 1 CBS
Output Poisoned sample set $U$, poisoned label set $S_{p}$.
  Initialize poisoned sample set $U=\{\}$
  for $i=1,...,N$ do
        $U=U\cup\{(x_{i},y_{i})\}$
     end if
  end for
  Return poisoned sample set $U$